Cygnus has a dividend cover ratio of 4.0 times and expects zero growth in dividends. The company has one million $1 ordinary shares in issue and the market capitalization (value) of the company is $50 million. After-tax profits for next year are expected to be $20 million.

What is the cost of equity capital?


A company has just declared dividend of GH¢50 per share and this is yet to be paid. The dividend is expected to grow at 10% per annum forever, and price cum Div is GH¢700 per share. What is the cost of equity?


A stock had an initial price of GH¢5,000 per share, paid a dividend of GH¢1,000 per share at the end of the year and had an ending share price of GH¢3,500.

What is (a)the dividend yield (b) the capital gain (%)  (c) total return on the stock



1
Expert's answer
2021-03-10T07:28:23-0500

1.


"50 billion+(20 billion-4.0\times 1billion)=66 billion"

2.


"700=50(1+0.1)^n"

"n=log_{1.1}14"

"n=27.6"

"50\times 27.6=1384.5"

3.

a)


"\frac{1000}{5000}\times 100=20"

b)


"\frac{3500-5000}{5000}=-0.3"

c)


"5000-(3500+1000)=-500"